### v1

# An alternative condition for the solvability of the Dirichlet problem for the minimal surface equation on non-mean convex domains

Abstract
We propose an alternative condition for the solvability of the Diri-chlet problem for the minimal surface equation that applies to non-mean convex domains. This condition is derived from a second-order ordinary differential equation whose solution produces a barrier that appears to be novel in the context of barrier constructions. It admits an explicit formulation and, in the setting of Hadamard manifolds, reveals a direct and transparent relationship between the geometry of the domain and the behavior of the boundary data required for solvability. The condition also extends naturally to unbounded domains.
In the Euclidean case, it is not only more practical to verify but also less restrictive than the classical Jenkins - Serrin criterion, ensuring the existence of solutions in situations where that approach fail. Furthermore, unlike the Jenkins-Serrin condition, our appproach separates the geometric properties of the domain from its boundary data, providing a clearer and more manageable framework for solvability analysis.
1 Introduction
Let $M$ be a complete $n$-dimensional Riemannian manifold, $n\geq 2$, and let $\Omega\varsubsetneq M$ be a $C^{2}$-domain. The graph of a function $u\in C^{2}(\Omega)$ is a minimal graph in $M\times\mathbb{R}$ if and only if $\mathcal{M}(u)=0$ in $\Omega$, where
$\mathcal{M}(u):=\left(1+|\nabla u|^{2}\right)\Delta u-\nabla^{2}u(\nabla u,\nabla u).$ (1)
We consider the Dirichlet problem:
$\mathcal{M}(u)$ $=0\quad\text{in }\Omega,\quad u\in C^{2}(\overline{\Omega}),$
$u|_{\partial\Omega}$ $=\varphi|_{\partial\Omega},$ (2)
where $\varphi\in C^{2}(\overline{\Omega})$ is given.
Let $H_{\partial\Omega}$ denote the mean curvature of $\partial\Omega$ with respect to the inner orientation, and define
$\partial\Omega^{-}:=\mathrm{clos}\left\{p\in\partial\Omega\,;\,H_{\partial\Omega}(p)<0\right\}.$
In the case of bounded domains, it is well known that if $\Omega\subset\mathbb{R}^{n}$, the Dirichlet problem for the minimal surface equation admits a solution for arbitrary continuous boundary data if and only if $\partial\Omega^{-}=\emptyset$ ([6], [8]). The existence part of this result has been generalized to Riemannian manifolds (see, e.g., [4]). If $\partial\Omega^{-}\neq\emptyset$, then, as shown in Theorem 1 of [3], which generalizes the classical result of H. Jenkins and J. Serrin (Theorem 2 of [8]), a bound on the oscillation of the boundary data $\varphi$ is sufficient to ensure solvability of (2), namely,
$\omega_{\varphi}(\partial\Omega):=\sup_{\partial\Omega}\varphi-\inf_{\partial\Omega}\varphi\leq B\left(|D\varphi|,|D^{2}\varphi|,|A|,\mathrm{Ric}_{M}\right),$ (3)
where $|A|$ is the norm of the second fundamental form of $\partial\Omega$, and $B$ is an explicit function (see [3], Section 2, p. 78 and Lemma 4; see also [8], Section 3, p. 179).
However, even in the Euclidean case, it is quite complicated to explicitly verify whether a given $C^{2}$ boundary function satisfies the oscillation bounds required in the above results, since the expression for $B$ depends on a neighborhood of each point in $\partial\Omega^{-}$, which is hard to describe explicitly. Furthermore, $B$ involves both the function and the domain, making it difficult to treat them separately. To illustrate this, we briefly explain the function $B$:
In [8], $B$ is constructed as follows. Since $\partial\Omega$ is compact and of class $C^{2}$, there exists $l>0$ (uniform for all $p\in\partial\Omega$) such that the intersection $B_{l}(p)\cap\partial\Omega$ consists of a single connected component that is the graph over a domain $D$ in a hyperplane parallel to $T_{p}\partial\Omega$ of a function $g\in C^{2}(D)$ satisfying $|\nabla g|<1$. Set $l^{\prime}=l/\sqrt{2}$. Also, since $\partial\Omega$ is compact and $C^{2}$, its principal curvatures are uniformly bounded; let $k$ be an absolute bound for these curvatures. Define
$\mathcal{A}$ $=\max\left\{\frac{\pi}{l^{\prime}},k,|D^{2}\varphi|\right\},$
$\mathcal{C}$ $=\frac{\mathcal{A}}{(1+|D\varphi|^{2})^{8n}},$
$\mathcal{H}$ $=\max_{\partial\Omega}(-H_{\partial\Omega}).$
Then $B$ is defined as
$B:=\frac{1}{16n\mathcal{A}}\cdot\begin{cases}\frac{\mathcal{C}}{\mathcal{H}},&\text{if }\mathcal{C}\leq\mathcal{H},\\ 1+\log\left(\frac{\mathcal{C}}{\mathcal{H}}\right),&\text{if }0<\mathcal{H}<\mathcal{C},\\ +\infty,&\text{if }\mathcal{H}\leq 0.\end{cases}$ (4)
In [3], $B$ is defined through an expression involving $|A|$ and a distance $\xi^{\prime}$ to $\partial\Omega$, $0<\xi^{\prime}\leq d_{0}$, being $B=+\infty$ at the mean convex points, where $d_{0}$ is such that the normal exponential map
$\exp:\partial\Omega\times[0,d_{0})\longrightarrow U_{d_{0}}=\{z\in\Omega\,;\,d(z)<d_{0}\}$ (5)
is a diffeomorphism (see Lemma 4 of [3]). Here, $\xi^{\prime}$ plays the same role as $l^{\prime}$ in the Jenkins-Serrin construction.
Given the above, we propose an alternative criterion to determine whether the Dirichlet problem (2) admits a solution for a given $\varphi\in C^{2}(\overline{\Omega})$. Our goal is to retain the geometric insights of previous results, while providing a more practical condition, at least in the Euclidean case, that also applies to unbounded domains and, in the case of Hadamard manifolds, can be phrased in a way that makes explicit the relationship between the geometry of the domain and the behavior of the boundary data that guarantees solvability (see Corollary 4). We are inspired by [10], which uses assumptions relating the $C^{2}$-norm of $\varphi$ and $\omega_{\varphi}(\overline{\Omega})$ with the radius $r$ of the circles that, at each $p\in\partial\Omega$, is tangent to $\partial\Omega$ at $p$ and lies in $\mathbb{R}^{2}\backslash\Omega$ (exterior circle condition). Our criterion arises from solving a second-order ODE (29), which appears to be a novel method for constructing barriers for the Dirichlet problem (2).
In special cases e.g., when $\Omega$ is bounded and the $C^{2}$ boundary data lies between two horizontal slices it is possible to provide simpler sufficient conditions involving $\omega_{\varphi}(\partial\Omega)$ to ensure solvability (see [1], [5]). Conditions on the oscillation are also used for less regular boundary data, but they are naturally more intricate (see [12], [9], [2], etc.).
In order to state our results, we introduce some notation. Let $i:M\rightarrow[0,\infty]$ denote the injectivity radius function of $M$.
Definition 1
We say that $\partial\Omega^{-}$ satisfies the exterior sphere condition of radius $r>0$ if $\partial\Omega^{-}\neq\emptyset$ and, at each $p\in\partial\Omega^{-}$, there exists a geodesic sphere $\Sigma_{p}\subset M\setminus\Omega$ of radius $r$ tangent to $\partial\Omega$ at $p$, with $r<R$, where
$R:=\inf\left\{i(p_{0})\,;\,p_{0}\text{ is the center of }\Sigma_{p},\,p\in\partial\Omega^{-}\right\}.$ (6)
This is abbreviated as the $r$-es condition.
Definition 2
We say that $\partial\Omega^{-}$ satisfies the locally Hadamard $r$-es condition if it satisfies the $r$-es condition and, for each $p\in\partial\Omega^{-}$, letting $\Sigma_{p,s}:=d_{p}^{-1}(s)$ for $s\in[0,R-r]$ (with $\Sigma_{p,0}=\Sigma_{p}$ and $d_{p}=d_{M}(\cdot,\Sigma_{p})$), and denoting $H_{\Sigma_{p,s}}$ the mean curvature of $\Sigma_{p,s}$ with respect to the outward orientation, we have:
$H_{\Sigma_{p}}(p)$ $\leq H_{\Sigma_{p,s}}(x),$ (7)
$|\nabla^{2}d_{p}|(x)$ $\leq\max_{\Sigma_{p}}|\nabla^{2}d_{p}|,$ (8)
for all $x\in\Sigma_{p,s}\cap\Omega$ and $s\in[0,R-r]$.
If $M$ is a Hadamard manifold and $\partial\Omega^{-}$ satisfies the $r$-es condition, then it satisfies the locally Hadamard $r$-es condition with $R=\infty$. The name reflects that the curvature behavior of the geodesic spheres $\Sigma_{p,s}$ in this case mimics that of Hadamard manifolds.
Let us define:
$\mu_{r}$ $:=\sup\left\{\max_{\Sigma_{p}}|\nabla^{2}d_{p}|\,;\,p\in\partial\Omega^{-}\right\},$ (9)
$\lambda_{r}$ $:=\inf\left\{H_{\Sigma_{p}}(p)\,;\,p\in\partial\Omega^{-}\right\}.$ (10)
Note that $\lambda_{r}<0$. For $M=\mathbb{R}^{n}$, we have $\lambda_{r}=-1/r$ and $\mu_{r}=1/r$.
Let
$\tau=\tau(\varphi):=\max\left\{\sup_{\overline{\Omega}}|\Delta\varphi|,\ \sup_{\overline{\Omega}}|\nabla\varphi|,\ \sup_{\overline{\Omega}}|\nabla^{2}\varphi|\right\},$ (11)
$\varrho:=\tau(1+2\tau^{2}),$ (12)
$a$ $:=\frac{2\tau^{3}+4\tau^{2}+3\tau}{\varrho},$ (13)
$b$ $:=-\frac{\tau^{2}+2\tau+2}{\varrho},$ (14)
$c$ $:=\frac{\tau^{2}}{\varrho},$ (15)
and denote $\omega:=\omega_{\varphi}(\overline{\Omega})$.
If $\partial\Omega^{-}$ satisfies the $r$-es condition, define
$\theta:=a+b(n-1)\lambda_{r}+c\mu_{r}.$ (16)
Note that $1<a<\theta$ (the maximum possible value of $a$ is approximately $3.7321$).
We now state the main result.
Theorem 3
Let $\Omega\subset M^{n}$ be a $C^{2}$-domain, where $M$ is a complete Riemannian manifold, and let $\varphi\in C^{2}(\overline{\Omega})$ be given. Assume that $\tau<\infty$ if $\Omega$ is unbounded, where $\tau=\tau(\varphi)$ is defined in (11). If either:
  • (i)
    $\partial\Omega^{-}=\emptyset$, or
  • (ii)
    $\partial\Omega^{-}$ satisfies the locally Hadamard $r$-es condition, and
    $\omega\leq\frac{1}{\varrho(\theta-1)^{2}}\left[\theta\left(1-e^{-\varrho(\theta-1)\delta}\right)-\varrho(\theta-1)\delta\right],$ (17)
    where $\varrho$ and $\theta$ are given by (12) and (16), respectively, and
    $\delta<\min\left\{R-r,\ \frac{\ln\theta}{\varrho(\theta-1)}\right\},$ (18)
then the Dirichlet problem (2) admits a solution. Moreover, the solution is unique if $\Omega$ is bounded.
Corollary 4
Let $\Omega\subset M^{n}$ be a $C^{2}$-domain, where $M$ is a Hadamard manifold, and let $\varphi\in C^{2}(\overline{\Omega})$ be given. Assume that $\tau<\infty$ if $\Omega$ is unbounded, where $\tau=\tau(\varphi)$ is defined in (11). If either:
  • (i)
    $\partial\Omega^{-}=\emptyset$, or
  • (ii)
    $\partial\Omega^{-}$ satisfies the $r$-es condition, and
    $\frac{\sqrt{a-\ln a-1}-\sqrt{\varrho\omega}(a+c\mu_{r}-1)}{\sqrt{\varrho\omega}\,b(n-1)}\leq\lambda_{r},$ (19)
then the Dirichlet problem (2) admits a solution. Moreover, the solution is unique if $\Omega$ is bounded.
We remark that on the right-hand side of inequality (19), we have only geometric information about the domain (specifically $\lambda_{r},$ the infimum of the mean curvature of the non mean convex part of the boundary of $\Omega$), while the left-hand side depends on the function $\varphi$ and the ambient manifold $M$ (via $\mu_{r}$).
In the particular case where $M=\mathbb{R}^{n}$ and $\partial\Omega^{-}$ satisfies the $r$-es condition, we have $\lambda_{r}=-1/r$ and $\mu_{r}=1/r$, so inequality (19) becomes:
$0<\frac{\sqrt{\varrho\omega}\left(c-b(n-1)\right)}{\sqrt{a-\ln a-1}-(a-1)\sqrt{\varrho\omega}}\leq r.$ (20)
Therefore, our condition is less restrictive than that of [10], where $r\geq 1$ is required and, moreover, holds for arbitrary dimension. It also allows us to conclude the solvability of the Dirichlet problem (2) in examples where the Jenkins-Serrin condition cannot be applied, as illustrated at the end of Section 3.
Regarding the connection with the Jenkins–Serrin condition, we emphasize that, even though our argument depends on an extension of $\varphi\in C^{2}(\partial\Omega)$ into $\overline{\Omega}$, it is always possible to construct such an extension - still denoted by $\varphi$ - on a neighborhood $\Omega_{\varepsilon}\subset\Omega$ of $\partial\Omega$, for some $\varepsilon>0$, in such a way that $\sup_{\overline{\Omega}_{\varepsilon}}\left|\nabla\varphi\right|=\sup_{\partial\Omega}\left|D\varphi\right|$ and $\sup_{\overline{\Omega}_{\varepsilon}}\left|\nabla^{2}\varphi\right|=\sup_{\partial\Omega}\left|D^{2}\varphi\right|$, where here $D\varphi$ and $D^{2}\varphi$ denote the first and second tangential derivatives of $\varphi$ on $\partial\Omega$. (See, for example, the extension constructed in Section 2.1 of [3].)
1.1 Basic context
Let $\left\langle\cdot,\cdot\right\rangle$ and $\nabla$ denote the Riemannian metric and the Riemannian connection of $M$, respectively. Let $\Omega\subset M$ be a $C^{2}$ domain and suppose that $\partial\Omega^{-}$ satisfies the $r$-es condition. Fix a point $p\in\partial\Omega^{-}$ and let $p_{0}$ be the center of the corresponding hypersurface $\Sigma_{p}$, as given in Definition 1. Define
$\Omega_{R}:=B_{R}(p_{0})\cap\Omega,$
where $R$ is given by (6). Consider the distance function
$d(x):=d_{p}(x)=d_{M}(x,\Sigma_{p})=d_{M}(x,p_{0})-r,\quad x\in\Omega_{R}.$
Since $\nabla d$ is orthogonal to the level sets $\Sigma_{p,s}:=d^{-1}(s)$, for $s=d(x)$, with $\Sigma_{p}=\Sigma_{0}$ and $s\in[0,R-r)$, we can define an orthonormal frame in a neighborhood of any given point $x\in\Omega_{R}$ as follows. First, near $x\in\Sigma_{p,s}\cap\Omega_{R}$, we choose an orthonormal frame $\{E_{i}\}_{i=1}^{n-1}$ tangent to $\Sigma_{p,s}$. We then extend this frame to a neighborhood $\Lambda\subset\Omega_{R}$ of $x$ by parallel transport along the geodesic rays emanating from $p_{0}$, and we continue to denote the extension by $\{E_{i}\}_{i=1}^{n-1}$. Define $E_{n}:=\nabla d$ on $\Lambda$. Then, $\{E_{i}\}_{i=1}^{n}$ forms an orthonormal frame on $\Lambda\subset\Omega_{R}$. [^1] Without loss of generality, we may assume from the outset that $\Lambda=B_{R}(p_{0})\cap\Omega$, i.e., $\Lambda=\Omega_{R}$.
Observe that, since the vector fields $E_{i}$ are parallel along the geodesics issuing from $p_{0}$, we have $\nabla_{E_{n}}E_{i}=0$ on $\Lambda$ for all $i=1,\dots,n$.
Lemma 5
In a neighborhood $\Lambda\subset\Omega_{R}$ as described above, we have:
$\nabla^{2}d(E_{i},E_{n})=0,\quad\text{for }i=1,\dots,n,$
and
$\nabla^{2}d(E_{i},E_{j})=-H_{E_{n}}(E_{i},E_{j}),\quad\text{for }i,j=1,\dots,n-1,$
where
$H_{E_{n}}(E_{i},E_{j}):=\left\langle-(\nabla_{E_{i}}E_{n})^{\top},E_{j}\right\rangle.$
In particular, $\nabla^{2}d(E_{i},E_{i})=-II_{E_{n}}(E_{i})$, where $II$ denotes the second fundamental form of the hypersurfaces $\Sigma_{p,s}$, and
$\Delta d=-\sum_{i=1}^{n-1}II_{E_{n}}(E_{i}).$
Proof. Straightforward calculation.
Lemma 6
Let $\varphi\in C^{2}(\overline{\Omega})$ and $\psi\in C^{\infty}([0,\infty))$ be given. Define the function $w\in C^{2}(\overline{\Lambda})$ by
$w=\varphi+\psi\circ d,$ (21)
where $\Lambda\subset\Omega_{R}$ is a neighborhood as defined above. Then,
$\nabla^{2}w(\nabla w,\nabla w)$ $=\left|\nabla\varphi\right|^{2}\nabla^{2}\varphi\left(\frac{\nabla\varphi}{\left|\nabla\varphi\right|},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)$
$\quad+\psi^{\prime}(d)\left[2\left|\nabla\varphi\right|\nabla^{2}\varphi\left(E_{n},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)+\left|\nabla\varphi\right|^{2}\nabla^{2}d\left(\frac{\nabla\varphi}{\left|\nabla\varphi\right|},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)\right]$
$\quad+\left[\psi^{\prime}(d)\right]^{2}\nabla^{2}\varphi(E_{n},E_{n})+\psi^{\prime\prime}(d)\left(E_{n}(\varphi)+\psi^{\prime}(d)\right)^{2}.$ (22)
Proof. A rather extensive computation yields
$\nabla_{\nabla w}\nabla w$ $=\nabla_{\nabla\varphi}\nabla\varphi+\psi^{\prime}(d)\left[\nabla_{E_{n}}\nabla\varphi+\nabla_{\nabla\varphi}E_{n}+\psi^{\prime\prime}(d)E_{n}\right]+\psi^{\prime\prime}(d)E_{n}(\varphi)E_{n}.$
Now, observing that $\left\langle\nabla_{\nabla\varphi}E_{n},E_{n}\right\rangle=0$, we conclude that
$\left\langle\nabla_{\nabla w}\nabla w,\nabla w\right\rangle$ $=\nabla^{2}\varphi(\nabla\varphi,\nabla\varphi)+\psi^{\prime}(d)\left[2\nabla^{2}\varphi(E_{n},\nabla\varphi)+\nabla^{2}d(\nabla\varphi,\nabla\varphi)\right]$
$\quad+\left[\psi^{\prime}(d)\right]^{2}\nabla^{2}\varphi(E_{n},E_{n})+\psi^{\prime\prime}(d)\left(E_{n}(\varphi)+\psi^{\prime}(d)\right)^{2},$
which proves the result.
Proposition 7
Assume that $\partial\Omega^{-}$ satisfies the locally Hadamard $r$-es condition. Let $\varphi\in C^{2}\left(\overline{\Omega}\right)$ be given, and let $\sigma:=\varrho\left(1-\theta\right)$, where $\varrho$ and $\theta$ are given by (12) and (16), respectively. Consider the function
$\psi(s)=\frac{1}{\sigma}\left[\varrho s+\left(\frac{\sigma-\varrho}{\sigma}\right)\left(e^{\sigma s}-1\right)\right],\quad s\in[0,\delta],$ (23)
where
$\delta<\min\left\{R-r,\frac{1}{\sigma}\ln\left(\frac{\varrho}{\varrho-\sigma}\right)\right\},$
and set $\Lambda_{p,\delta}=B_{r+\delta}(p_{0})\cap\Omega$. Then the function $w=\varphi+\psi\circ d$ belongs to $C^{2}(\overline{\Lambda}_{p,\delta})$ and satisfies
$w(p)=\varphi(p),\quad w(x)\geq\varphi(x),\quad\text{and}\quad\mathcal{M}(w(x))\leq 0\quad\text{for all }x\in\overline{\Lambda}_{p,\delta}.$
Proof. We explain how the function (23) was constructed. We start by assuming only that $\psi\in C^{\infty}\left([0,\infty\right)$ and satisfies the conditions $\psi\left(0\right)=0$, $\psi^{\prime}\left(s\right)>0$ and $\psi^{\prime\prime}\left(s\right)<0$ for $s>0$, necessary conditions for $\psi$ to serve as an upper barrier. We assume, from now onward, that $\Lambda\subset\Omega_{R}$ is such that $p\in\partial\Lambda$. From (1), we have $\mathcal{M}(w)(x)\leq 0$ in $\Lambda$ if
$(1+|\nabla w|^{2})\Delta w(x)-\nabla^{2}w(\nabla w,\nabla w)(x)\leq 0,$
where $w=\varphi+\psi\circ d$. By Lemma 6, $\nabla^{2}w(\nabla w,\nabla w)$ can be expressed as in (22). Since
$\left(1+\left|\nabla w\right|^{2}\right)\Delta w=\left(1+\left|\nabla w\right|^{2}\right)\Delta\varphi+(1+\left|\nabla w\right|^{2})\psi^{\prime\prime}\left(d\right)+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d,$
it follows that $\mathcal{M}(w)(x)\leq 0$ if the following sum is nonpositive:
$\left(1+\left|\nabla w\right|^{2}\right)\Delta\varphi+(1+\left|\nabla w\right|^{2})\psi^{\prime\prime}\left(d\right)+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d+$
$-\left|\nabla\varphi\right|^{2}\nabla^{2}\varphi\left(\frac{\nabla\varphi}{\left|\nabla\varphi\right|},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)+$
$-2\psi^{\prime}\left(d\right)\left|\nabla\varphi\right|\nabla^{2}\varphi\left(E_{n},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)-\psi^{\prime}\left(d\right)\left|\nabla\varphi\right|^{2}\nabla^{2}d\left(\frac{\nabla\varphi}{\left|\nabla\varphi\right|},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)+$
$-\left[\psi^{\prime}\left(d\right)\right]^{2}\nabla^{2}\varphi(E_{n},E_{n})-\psi^{\prime\prime}\left(d\right)\left(E_{n}\left(\varphi\right)+\psi^{\prime}\left(d\right)\right)^{2}.$
In $\Lambda$, we estimate:
$-|\nabla\varphi|^{2}\nabla^{2}\varphi\left(\frac{\nabla\varphi}{|\nabla\varphi|},\frac{\nabla\varphi}{|\nabla\varphi|}\right)$ $\leq\tau^{3},$
$-|\nabla\varphi|\nabla^{2}\varphi\left(E_{n},\frac{\nabla\varphi}{|\nabla\varphi|}\right)$ $\leq\tau^{2},$
$-|\nabla\varphi|^{2}\nabla^{2}d\left(\frac{\nabla\varphi}{|\nabla\varphi|},\frac{\nabla\varphi}{|\nabla\varphi|}\right)$ $\leq\tau^{2}\mu_{r},$
$-\nabla^{2}\varphi(E_{n},E_{n})$ $\leq\tau.$
where $\tau$ is given by (11). The third inequality uses that $\nabla^{2}d_{p}|_{\Sigma_{p,s}}\leq\mu_{r}$ for all $s\in[0,R]$, due to the locally Hadamard $r$-es condition. Therefore, the sum is nonpositive in $\Lambda$ if
$\left(1+\left|\nabla w\right|^{2}\right)\tau+(1+\left|\nabla w\right|^{2})\psi^{\prime\prime}\left(d\right)+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d+$
$+\tau^{3}+2\psi^{\prime}\left(d\right)\tau^{2}+\psi^{\prime}\left(d\right)\tau^{2}\mu_{r}$
$+\left[\psi^{\prime}\left(d\right)\right]^{2}\tau-\psi^{\prime\prime}\left(d\right)\left(E_{n}\left(\varphi\right)+\psi^{\prime}\left(d\right)\right)^{2}.$ (24)
Now compute
$\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime\prime}\left(d\right)$ $=$ $\psi^{\prime\prime}\left(d\right)+\left|\nabla\varphi\right|^{2}\psi^{\prime\prime}\left(d\right)+$
$+2\psi^{\prime}\left(d\right)E_{n}\left(\varphi\right)\psi^{\prime\prime}\left(d\right)+\left[\psi^{\prime}\left(d\right)\right]^{2}\psi^{\prime\prime}\left(d\right)$
and, since $\psi^{\prime\prime}<0$, we obtain
$\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime\prime}\left(d\right)-\psi^{\prime\prime}\left(d\right)\left(E_{n}\left(\varphi\right)+\psi^{\prime}\left(d\right)\right)^{2}\leq\psi^{\prime\prime}\left(d\right).$
Then (24) is non positive in $\Lambda$ if
$\left(1+\left|\nabla w\right|^{2}\right)\tau+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d+$
$\tau^{3}+2\psi^{\prime}\left(d\right)\tau^{2}+\psi^{\prime}\left(d\right)\tau^{2}\mu_{r}$
$+\left[\psi^{\prime}\left(d\right)\right]^{2}\tau+\psi^{\prime\prime}\left(d\right)\leq 0.$ (25)
We have
$\left(1+\left|\nabla w\right|^{2}\right)\tau$ $\leq$ $\left(1+\left|\nabla\varphi\right|^{2}+2\psi^{\prime}\left(d\right)\left|\nabla\varphi\right|+\left[\psi^{\prime}\left(d\right)\right]^{2}\right)\tau$
$\leq$ $\tau+\tau^{3}+2\tau^{2}\psi^{\prime}\left(d\right)+\tau\left[\psi^{\prime}\left(d\right)\right]^{2}.$
Thus, inequality (25) holds if
$\tau+2\tau^{3}+4\tau^{2}\psi^{\prime}\left(d\right)+2\tau\left[\psi^{\prime}\left(d\right)\right]^{2}+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d+$
$+\psi^{\prime}\left(d\right)\tau^{2}\mu_{r}+\psi^{\prime\prime}\left(d\right)$ (26)
is nonpositive. Using Lemma 5, for $x\in\Lambda$ and $s=d(x)\in[0,R)$, we have
$\Delta d(x)=-\sum_{i=1}^{n-1}II_{E_{n}}(E_{i})(x)=-(n-1)H_{\Sigma_{p,s}}(x),$
where $H_{\Sigma_{p,s}}$ is the mean curvature of the immersion $\Sigma_{p,s}$ with respect to $E_{n}=\nabla d$. Since $p\in\partial\Omega^{-}$, we have
$H_{\Sigma_{p}}(p)\leq H_{\partial\Omega}(p)<0,$
and from the locally Hadamard $r$-es condition,
$\lambda_{r}\leq H_{\Sigma_{p}}(p)\leq H_{\Sigma_{p,s}}(x),$
for all $x\in\Sigma_{p,s}\cap\Omega_{R}$ and $s\in[0,R-r]$. Hence,
$\Delta d(x)\leq-(n-1)\lambda_{r}$
and we have (26) if
$\tau+2\tau^{3}+4\tau^{2}\psi^{\prime}\left(d\right)+2\tau\left[\psi^{\prime}\left(d\right)\right]^{2}-\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\left(n-1\right)\lambda_{r}+$
$+\psi^{\prime}\left(d\right)\tau^{2}\mu_{r}+\psi^{\prime\prime}\left(d\right)\leq 0.$ (27)
Observing that
$\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\leq\left(1+\tau^{2}\right)\psi^{\prime}\left(d\right)+2\left[\psi^{\prime}\left(d\right)\right]^{2}\tau\mathfrak{+}\left[\psi^{\prime}\left(d\right)\right]^{3}$
it follows that we have (27) in $\Lambda$ if
$\psi^{\prime\prime}\left(d\right)$ $\leq$ $\left(n-1\right)\lambda_{r}\left[\psi^{\prime}\left(d\right)\right]^{3}-2\tau\left[1-\left(n-1\right)\lambda_{r}\right]\left[\psi^{\prime}\left(d\right)\right]^{2}$
$-\left[\left(4+\mu_{r}\right)\tau^{2}-\left(1+\tau^{2}\right)\left(n-1\right)\lambda_{r}\right]\psi^{\prime}\left(d\right)$
$-\tau\left[1+2\tau^{2}\right].$
Letting
$\alpha:=2\tau[1-(n-1)\lambda_{r}],\quad\beta:=(4+\mu_{r})\tau^{2}-(1+\tau^{2})(n-1)\lambda_{r},$
we write the inequality above as
$\psi^{\prime\prime}\leq\left(n-1\right)\lambda_{r}\left[\psi^{\prime}\right]^{3}-\alpha\left[\psi^{\prime}\right]^{2}-\beta\psi^{\prime}-\varrho.$ (28)
Assume $0<\psi^{\prime}\leq 1$. Then
$\left(n-1\right)\lambda_{r}\psi^{\prime}$ $\leq$ $\left(n-1\right)\lambda_{r}\left[\psi^{\prime}\right]^{3}<0$
$-\alpha\psi^{\prime}$ $\leq$ $-\alpha\left[\psi^{\prime}\right]^{2}<0$
so inequality (28) is satisfied if $\psi^{\prime\prime}-\sigma\psi^{\prime}+\varrho\leq 0$, with $\sigma=\varrho(1-\theta)<0$ since $\theta>1$. Thus, we solve the ODE
$\psi^{\prime\prime}-\sigma\psi^{\prime}+\varrho=0,$ (29)
with initial condition $\psi(0)=0$, whose solution is
$\psi(s)=\frac{1}{\sigma}\left[\varrho s-c_{1}(e^{\sigma s}-1)\right].$
We require $0<\psi^{\prime}(s)\leq 1$, i.e.,
$0<\frac{1}{\sigma}(\varrho-c_{1}\sigma e^{\sigma s})\leq 1,$
which implies
$\frac{\varrho}{\sigma}e^{-\sigma s}<-c_{1}\leq\left(\frac{\sigma-\varrho}{\sigma}\right)e^{-\sigma s}.$
Let $\rho=R-r$. Then everything is in order if
$\frac{\varrho-\sigma}{\sigma}\leq c_{1}<\frac{\varrho}{\sigma}e^{-\sigma\rho}.$ (30)
This holds if
$\ln\left(\frac{\varrho}{\varrho-\sigma}\right)<\sigma\rho.$ (31)
As $\frac{\varrho}{\varrho-\sigma}<1$, condition (31) is satisfied if
$\rho<\frac{1}{\sigma}\ln\left(\frac{\varrho}{\varrho-\sigma}\right).$
Therefore, if
$\delta<\min\left\{R-r,\frac{1}{\sigma}\ln\left(\frac{\varrho}{\varrho-\sigma}\right)\right\},$ (32)
then, for
$\psi(s)=\frac{1}{\sigma}\left[\varrho s-c_{1}(e^{\sigma s}-1)\right],\quad s\in[0,\delta],$
with $c_{1}=\frac{\varrho-\sigma}{\sigma}$, we have $\psi(0)=0$, $0<\psi^{\prime}(s)\leq 1$, and $\psi^{\prime\prime}(s)=-c_{1}\sigma e^{\sigma s}<0$. Furthermore, $\psi(s)\geq 0$ on $0,\delta]$. Finally, defining
$\Lambda_{p,\delta}=B_{r+\delta}(p_{0})\cap\Omega,$
we obtain the function
$w(x)=\varphi(x)+\psi(d(x)),$
with
$\psi(s)=\frac{1}{\sigma}\left[\varrho s+\left(\frac{\sigma-\varrho}{\sigma}\right)(e^{\sigma s}-1)\right],\quad 0\leq s\leq\delta,$ (33)
which satisfies $w(p)=\varphi(p)$, $w(x)\geq\varphi(x)$ for all $x\in\overline{\Lambda}_{p,\delta}$, and $\mathcal{M}(w(x))\leq 0$ in $\overline{\Lambda}_{p,\delta}$.
1.2 Proofs of Theorem 3 and Corollary 4
Proof of Theorem 3.
In the case where $\partial\Omega^{-}=\emptyset$ and $\Omega$ is bounded, the result follows, for instance, from Theorem 1 of [3].
Suppose $\Omega$ bounded and that $\partial\Omega^{-}$ satisfies the locally Hadamard $r$-es condition.
Given any $p\in\partial\Omega^{-}$, consider the neighborhood $\Lambda_{p,\delta}$ as defined in Proposition 7, where $\delta$ satisfies condition (32). Define the function
$w_{p}=\varphi+\psi(d_{p}),$
where $\psi$ is as in (33). To construct a barrier from above at $p$, we require that
$w_{p}(x)=\varphi(x)+\psi(\delta)\geq\sup_{\overline{\Omega}}\varphi\quad\text{on }\partial\Lambda_{p,\delta}\setminus\partial\Omega,$
which holds if
$\psi(\delta)\geq\sup_{\overline{\Omega}}\varphi-\inf_{\overline{\Omega}}\varphi=:\omega.$
Since $\delta$ is independent of the point $p\in\partial\Omega^{-}$, Proposition 7 implies that, if
$\omega\leq\frac{1}{\sigma}\left[\varrho\delta+\left(\frac{\sigma-\varrho}{\sigma}\right)\left(e^{\sigma\delta}-1\right)\right],$ (34)
then we can construct an upper barrier at each $p\in\partial\Omega^{-}$ by setting
$v_{p}(x)=\begin{cases}w_{p}(x),&\text{if }x\in\Lambda_{p,\delta},\\ \displaystyle\inf_{\partial\Lambda_{p,\delta}\setminus\partial\Omega}w_{p},&\text{if }x\in\Omega\setminus\Lambda_{p,\delta}.\end{cases}$
For a lower barrier at $p\in\partial\Omega^{-}$, we define $\xi_{p}:=\varphi-\psi\circ d$.
At points of $\partial\Omega$ where the domain is mean convex, we proceed as in the proof of Theorem 1 of [3], and no additional geometric assumptions are required to construct barriers.
The conclusion now follows from classical elliptic PDE theory [7].
If $\Omega$ is unbounded and $\tau<\infty$, the result follows directly from the method of Perron (see Theorem 5 of [11]). Indeed, define, at given $x\in\overline{\Omega},$
$u\left(x\right)=\sup\left\{s\left(x\right);s\in C^{0}\left(\Omega\right)\text{ is generalized solution of }M\text{ and }s|_{\partial\Omega}\leq\varphi|_{\partial\Omega}\right\}.$
Since every subsolution used to define $u$ is bounded above by $\sup_{\Omega}\varphi$, the function $u$ is well-defined. Moreover, by Perron’s method, we have $u\in C^{\infty}\left(\Omega\right)$ and $M\left(u\right)=0.$ The barriers at the boundary ensures that $u\in C^{2}\left(\overline{\Omega}\right)$ and $u|_{\partial\Omega}=\varphi.$
Proof of Corollary 4.
When $M$ is a Hadamard manifold, we have $R=\infty$, so we can take $\delta$ in (32) as
$\delta=\frac{1}{\sigma}\ln\left(\frac{\varrho}{\varrho-\sigma}\right).$
Substituting this into (34) and recalling that $\sigma=\varrho(1-\theta)$, condition (34) becomes
$\omega\leq\frac{1}{\varrho(\theta-1)^{2}}\left(\theta-\ln\theta-1\right).$ (35)
Since $1<a<\theta$ and observing that the function $f(x)=x-\ln x-1$ is positive and increasing for $x>1$, we have that
$\omega\leq\frac{a-\ln a-1}{\varrho(\theta-1)^{2}}$ (36)
implies (35). Now observe that
$\theta=a+b(n-1)\lambda_{r}+c\mu_{r}.$
Plugging this into (36) leads to the equivalent inequality
$\frac{\sqrt{a-\ln a-1}-\sqrt{\varrho\omega}(a+c\mu_{r}-1)}{\sqrt{\varrho\omega}\,b(n-1)}\leq\lambda_{r}.$ (37)
Remark 8
Since $\lambda_{r}<0$, the left-hand side of (37) must be negative. This occurs if
$\omega<\frac{a-\ln a-1}{\varrho(a+c\mu_{r}-1)^{2}},$
which is guaranteed by condition (36).
Example 9
We now present an example illustrating that our condition ensures the existence of a solution to the Dirichlet problem (2) even in cases where the Jenkins-Serrin criterion fails to do so.
Let $\Omega\subset\mathbb{R}^{2}$ be the bounded domain whose boundary is parametrized by
$\gamma(t)=\xi(t)(\cos t,\sin t),\quad t\in[0,2\pi],$
where
$\xi(t)=\sqrt{4\cos(2t)+\sqrt{16\cos^{2}(2t)+12}}.$
Let $\varphi\in C^{2}(\overline{\Omega})$ be given by
$\varphi(x,y)=\frac{e^{-y}}{20}+1.$
Figure 1: The domain used in the example
Regarding the geometry of $\partial\Omega$, we compute
$0.45\approx\lambda_{r}\leq k_{\partial\Omega}\leq k:=\sup_{\partial\Omega}|k_{\partial\Omega}|\approx 0.82,$
so that $\mu_{r}=-\lambda_{r}$. For the boundary data $\varphi|_{\partial\Omega}$, we find
$\omega_{\varphi}(\overline{\Omega})=\omega_{\varphi}(\partial\Omega)\approx 0.17,$
with $\tau\approx 0.1871$, $\theta\approx 9.05$, and the maximal admissible value of $\delta$ from (32) is approximately $1.37$. Choosing $\delta=1$ for simplicity, the right-hand side of inequality (17) becomes approximately $0.43$. Since $\omega\approx 0.17<0.43$, the condition is satisfied.
On the other hand, following the Jenkins-Serrin method, we proceed as in case I) above for computing $B$. Using the corresponding notation and in order to obtain the maximum value for $B$, we estimate:
$l=0.25,\quad l^{\prime}\approx 0.1767,\quad\frac{\pi}{l^{\prime}}\approx 17.7792,\quad k\approx 0.82,$
$|D^{2}\varphi|\approx 0.1871,\quad A\approx 17.7792,\quad C\approx 10.2526,\quad\mathcal{H}\approx 0.45.$
These values yield
$B\approx 0.0072<\omega_{\varphi}(\partial\Omega),$
so the Jenkins-Serrin condition fails, while our condition still guarantees existence.