### v3

# An alternative solvability criterion for the Dirichlet problem for the minimal surface equation and an application to the mean curvature flow

Abstract
We propose an alternative condition for the solvability of the Dirichlet problem for the minimal surface equation that applies to non-mean convex domains. This condition is derived from a second-order ordinary differential equation whose solution produces a barrier that appears to be novel in the context of barrier constructions. It admits an explicit formulation and, in the setting of Hadamard manifolds, reveals a direct and transparent relationship between the geometry of the domain and the behavior of the boundary data required for solvability. The condition also extends naturally to unbounded domains.
In the Euclidean case, it is not only more practical to verify but also less restrictive than the classical Jenkins–Serrin criterion, ensuring the existence of solutions in situations where that approach fails. Furthermore, unlike the Jenkins–Serrin condition, our approach separates the domain’s geometric properties from its boundary data, providing a clearer and more manageable framework for solvability analysis.
Beyond the stationary theory, the same ODE-generated barrier also yields natural sub- and supersolutions for the mean curvature flow of graphs, allowing one to initiate and control the short-time evolution from non-mean convex boundaries. As a consequence, our barrier construction extends a classical result of Ecker–Huisken on short-time existence of graphical mean curvature flow to domains with non-mean convex boundary; in the mean convex setting, their original result is recovered.
1 Introduction
Let $M$ be a complete $n$-dimensional Riemannian manifold, $n\geq 2$, and let $\Omega\varsubsetneq M$ be a $C^{2}$-domain. The graph of a function $u\in C^{2}(\Omega)$ is a minimal graph in $M\times\mathbb{R}$ if and only if $\mathcal{M}(u)=0$ in $\Omega$, where
$\mathcal{M}(u):=\left(1+|\nabla u|^{2}\right)\Delta u-\nabla^{2}u(\nabla u,\nabla u).$ (1)
We consider the Dirichlet problem:
$\mathcal{M}(u)$ $=0\quad\text{in }\Omega,\quad u\in C^{2}(\overline{\Omega}),$
$u|_{\partial\Omega}$ $=\varphi|_{\partial\Omega},$ (2)
where $\varphi\in C^{2}(\overline{\Omega})$ is given.
Let $H_{\partial\Omega}$ denote the mean curvature of $\partial\Omega$ with respect to the inner orientation, and define
$\partial\Omega^{-}:=\mathrm{clos}\left\{p\in\partial\Omega\,;\,H_{\partial\Omega}(p)<0\right\}.$
In the case of bounded domains, it is well known that if $\Omega\subset\mathbb{R}^{n}$, the Dirichlet problem for the minimal surface equation admits a solution for arbitrary continuous boundary data if and only if $\partial\Omega^{-}=\emptyset$ ([9], [11]). The existence part of this result has been generalized to Riemannian manifolds (see, e.g., [5]). If $\partial\Omega^{-}\neq\emptyset$, then, as shown in Theorem 1 of [3], which generalizes the classical result of H. Jenkins and J. Serrin (Theorem 2 of [11]), a bound on the oscillation of the boundary data $\varphi$ is sufficient to ensure solvability of (2), namely,
$\omega_{\varphi}(\partial\Omega):=\sup_{\partial\Omega}\varphi-\inf_{\partial\Omega}\varphi\leq B\left(|D\varphi|,|D^{2}\varphi|,|A|,\mathrm{Ric}_{M}\right),$ (3)
where $|A|$ is the norm of the second fundamental form of $\partial\Omega$, and $B$ is an explicit function (see [3], Section 2, p. 78 and Lemma 4; see also [11], Section 3, p. 179).
However, even in the Euclidean case, it is quite complicated to explicitly verify whether a given $C^{2}$ boundary function satisfies the oscillation bounds required in the above results, since the expression for $B$ depends on a neighborhood of each point in $\partial\Omega^{-}$, which is hard to describe explicitly. Furthermore, $B$ involves both the function and the domain, making it difficult to treat them separately. To illustrate this, we briefly explain the function $B$:
In [11], $B$ is constructed as follows. Since $\partial\Omega$ is compact and of class $C^{2}$, there exists $l>0$ (uniform for all $p\in\partial\Omega$) such that the intersection $B_{l}(p)\cap\partial\Omega$ consists of a single connected component that is the graph over a domain $D$ in a hyperplane parallel to $T_{p}\partial\Omega$ of a function $g\in C^{2}(D)$ satisfying $|\nabla g|<1$. Set $l^{\prime}=l/\sqrt{2}$. Also, since $\partial\Omega$ is compact and $C^{2}$, its principal curvatures are uniformly bounded; let $k$ be an upper bound. Define
$\mathcal{A}$ $=\max\left\{\frac{\pi}{l^{\prime}},k,|D^{2}\varphi|\right\},$
$\mathcal{C}$ $=\frac{\mathcal{A}}{(1+|D\varphi|^{2})^{8n}},$
$\mathcal{H}$ $=\max_{\partial\Omega}(-H_{\partial\Omega}).$
Then $B$ is defined as
$B:=\frac{1}{16n\mathcal{A}}\cdot\begin{cases}\frac{\mathcal{C}}{\mathcal{H}},&\text{if }\mathcal{C}\leq\mathcal{H},\\ 1+\log\left(\frac{\mathcal{C}}{\mathcal{H}}\right),&\text{if }0<\mathcal{H}<\mathcal{C},\\ +\infty,&\text{if }\mathcal{H}\leq 0.\end{cases}$ (4)
In [3], $B$ is defined through an expression involving $|A|$ and a distance $\xi^{\prime}\in(0,d_{0}]$, where $d_{0}$ is such that the normal exponential map
$\exp:\partial\Omega\times[0,d_{0})\longrightarrow U_{d_{0}}=\{z\in\Omega\,;\,d(z)<d_{0}\}$ (5)
is a diffeomorphism (see Lemma 4 of [3]). Here, $\xi^{\prime}$ plays the same role as $l^{\prime}$ in the Jenkins-Serrin construction.
Given the above, we propose an alternative criterion to determine whether the Dirichlet problem (2) admits a solution for a given $\varphi\in C^{2}(\overline{\Omega})$. Our goal is to retain the geometric insights of previous results, while providing a more practical condition at least in the Euclidean case that also applies to unbounded domains and, in the case of Hadamard manifolds, can be phrased in a way that makes explicit the relationship between the geometry of the domain and the behavior of the boundary data that guarantees solvability (see Corollary 4). We are inspired by [14], which uses assumptions relating the $C^{2}$-norm of $\varphi$ and $\omega_{\varphi}(\overline{\Omega})$ with the maximal radius of exterior tangent circles to $\partial\Omega$. Our criterion arises from solving a second-order ODE (29), which appears to be a novel method for constructing barriers for the Dirichlet problem (2).
In special cases e.g., when $\Omega$ is bounded and the $C^{2}$ boundary data lies between two horizontal slices it is possible to provide simpler sufficient conditions involving $\omega_{\varphi}(\partial\Omega)$ to ensure solvability (see [1], [8]). Conditions on the oscillation are also used for less regular boundary data, but they are naturally more intricate (see [16], [12], [2], etc.).
The ideas developed here extend naturally to the parabolic setting of graphical mean curvature flow. Consider the evolution
$\partial_{t}u=\mathcal{M}(u),$
where $\mathcal{M}$ is given by (1). The same differential inequality underlying the ODE that generates our barrier reappears in the derivation of a priori height and gradient controls for the parabolic problem. As a consequence, the structural conditions ensuring solvability of the stationary Dirichlet problem provide, at the same time, sufficient hypotheses to preserve graphicality and boundary regularity for short time along the flow, even when $\partial\Omega$ is not mean convex (see Section 2). In particular, our criterion yields a short-time preservation mechanism for graphical mean curvature flow that extends Theorem 1.1 of Ecker-Huisken [6] to the non-mean-convex setting, while recovering their result in the mean-convex case. This yields a bridge between the elliptic theory and the mean curvature flow of graphs and furnishes explicit sub/supersolutions usable at the parabolic level (see, e.g., [6, 4]).
In order to state our next results, we introduce some notation. Let $i:\nolinebreak M\rightarrow[0,\infty]$ denote the injectivity radius function of $M$.
Definition 1
We say that $\partial\Omega^{-}$ satisfies the exterior sphere condition of radius $r>0$ if $\partial\Omega^{-}\neq\emptyset$ and, at each $p\in\partial\Omega^{-}$, there exists a geodesic sphere $\Sigma_{p}\subset M\setminus\Omega$ of radius $r$ tangent to $\partial\Omega$ at $p$, with $r<R$, where
$R:=\inf\left\{i(p_{0})\,;\,p_{0}\text{ is the center of }\Sigma_{p},\,p\in\partial\Omega^{-}\right\}.$ (6)
This is abbreviated as the $r$-es condition.
Definition 2
We say that $\partial\Omega^{-}$ satisfies the locally Hadamard $r$-es condition if it satisfies the $r$-es condition and, for each $p\in\partial\Omega^{-}$, letting $\Sigma_{p,s}:=d_{p}^{-1}(s)$ for $s\in[0,R-r]$ (with $\Sigma_{p,0}=\Sigma_{p}$ and $d_{p}=d_{M}(\cdot,\Sigma_{p})$), and denoting $H_{\Sigma_{p,s}}$ the mean curvature of $\Sigma_{p,s}$ with respect to the outward orientation, we have:
$H_{\Sigma_{p}}(p)$ $\leq H_{\Sigma_{p,s}}(x),$ (7)
$|\nabla^{2}d_{p}|(x)$ $\leq\max_{\Sigma_{p}}|\nabla^{2}d_{p}|,$ (8)
for all $x\in\Sigma_{p,s}\cap\Omega$ and $s\in[0,R-r]$.
If $M$ is a Hadamard manifold and $\partial\Omega^{-}$ satisfies the $r$-es condition, then it satisfies the locally Hadamard $r$-es condition with $R=\infty$. The name reflects that the curvature behavior of the geodesic spheres $\Sigma_{p,s}$ in this case mimics that of Hadamard manifolds.
Let us define:
$\mu_{r}$ $:=\sup\left\{\max_{\Sigma_{p}}|\nabla^{2}d_{p}|\,;\,p\in\partial\Omega^{-}\right\},$ (9)
$\lambda_{r}$ $:=\inf\left\{H_{\Sigma_{p}}(p)\,;\,p\in\partial\Omega^{-}\right\}.$ (10)
Note that $\lambda_{r}<0$. For $M=\mathbb{R}^{n}$, we have $\lambda_{r}=-1/r$ and $\mu_{r}=1/r$.
Let
$\tau=\tau(\varphi):=\max\left\{\sup_{\overline{\Omega}}|\Delta\varphi|,\ \sup_{\overline{\Omega}}|\nabla\varphi|,\ \sup_{\overline{\Omega}}|\nabla^{2}\varphi|\right\},$ (11)
$\varrho:=\tau(1+2\tau^{2}),$ (12)
$a$ $:=\frac{2\tau^{3}+4\tau^{2}+3\tau}{\varrho},$ (13)
$b$ $:=-\frac{\tau^{2}+2\tau+2}{\varrho},$ (14)
$c$ $:=\frac{\tau^{2}}{\varrho},$ (15)
and denote $\omega:=\omega_{\varphi}(\overline{\Omega})$.
If $\partial\Omega^{-}$ satisfies the $r$-es condition, define
$\theta:=a+b(n-1)\lambda_{r}+c\mu_{r}.$ (16)
Note that $1<a<\theta$ (the maximum possible value of $a$ is approximately $3.7321$).
We now state the main result.
Theorem 3
Let $\Omega\subset M^{n}$ be a $C^{2}$-domain, where $M$ is a complete Riemannian manifold, and let $\varphi\in C^{2}(\overline{\Omega})$ be given. Assume that $\tau<\infty$ if $\Omega$ is unbounded, where $\tau=\tau(\varphi)$ is defined in (11). If either:
  • (i)
    $\partial\Omega^{-}=\emptyset$, or
  • (ii)
    $\partial\Omega^{-}$ satisfies the locally Hadamard $r$-es condition, and
    $\omega\leq\frac{1}{\varrho(\theta-1)^{2}}\left[\theta\left(1-e^{-\varrho(\theta-1)\delta}\right)-\varrho(\theta-1)\delta\right],$ (17)
    where $\varrho$ and $\theta$ are given by (12) and (16), respectively, and
    $\delta<\min\left\{R-r,\ \frac{\ln\theta}{\varrho(\theta-1)}\right\},$ (18)
then the Dirichlet problem (2) admits a solution. Moreover, the solution is unique if $\Omega$ is bounded.
Corollary 4
Let $\Omega\subset M^{n}$ be a $C^{2}$-domain, where $M$ is a Hadamard manifold, and let $\varphi\in C^{2}(\overline{\Omega})$ be given. Assume that $\tau<\infty$ if $\Omega$ is unbounded, where $\tau=\tau(\varphi)$ is defined in (11). If either:
  • (i)
    $\partial\Omega^{-}=\emptyset$, or
  • (ii)
    $\partial\Omega^{-}$ satisfies the $r$-es condition, and
    $\frac{\sqrt{a-\ln a-1}-\sqrt{\varrho\omega}(a+c\mu_{r}-1)}{\sqrt{\varrho\omega}\,b(n-1)}\leq\lambda_{r},$ (19)
then the Dirichlet problem (2) admits a solution. Moreover, the solution is unique if $\Omega$ is bounded.
We remark that on the right-hand side of inequality (19), we have only geometric information about the domain (specifically $\lambda_{r},$ the infimum of the mean curvature of the non mean convex part of the boundary of $\Omega$), while the left-hand side depends on the function $\varphi$ and the ambient manifold $M$ (via $\mu_{r}$).
In the particular case where $M=\mathbb{R}^{n}$ and $\partial\Omega^{-}$ satisfies the $r$-es condition, we have $\lambda_{r}=-1/r$ and $\mu_{r}=1/r$, so inequality (19) becomes:
$0<\frac{\sqrt{\varrho\omega}\left(c-b(n-1)\right)}{\sqrt{a-\ln a-1}-(a-1)\sqrt{\varrho\omega}}\leq r.$ (20)
Therefore, our condition is less restrictive than that of [14], where $r\geq 1$ is required and, moreover, holds for arbitrary dimension. It also allows us to conclude the solvability of the Dirichlet problem (2) in examples where the Jenkins-Serrin condition cannot be applied, as illustrated at the end of Section 1.2.
Regarding the connection with the Jenkins–Serrin condition, we emphasize that, even though our argument depends on an extension of $\varphi\in C^{2}(\partial\Omega)$ into $\overline{\Omega}$, it is always possible to construct such an extension—still denoted by $\varphi$—on a neighborhood $\Omega_{\varepsilon}\subset\Omega$ of $\partial\Omega$, for some $\varepsilon>0$, in such a way that
$\sup_{\overline{\Omega}_{\varepsilon}}|\nabla\varphi|=\sup_{\partial\Omega}|D\varphi|\quad\text{and}\quad\sup_{\overline{\Omega}_{\varepsilon}}|\nabla^{2}\varphi|=\sup_{\partial\Omega}|D^{2}\varphi|,$
where $D\varphi$ and $D^{2}\varphi$ denote the first and second tangential derivatives of $\varphi$ on $\partial\Omega$ (see, for example, the extension constructed in Section 2.1 of [3]).
1.1 Basic context
Let $\left\langle\cdot,\cdot\right\rangle$ and $\nabla$ denote the Riemannian metric and the Riemannian connection of $M$, respectively. Let $\Omega\subset M$ be a $C^{2}$ domain and suppose that $\partial\Omega^{-}$ satisfies the $r$-es condition. Fix a point $p\in\partial\Omega^{-}$ and let $p_{0}$ be the center of the corresponding hypersurface $\Sigma_{p}$, as given in Definition 1. Define
$\Omega_{R}:=B_{R}(p_{0})\cap\Omega,$
where $R$ is given by (6). Consider the distance function
$d(x):=d_{p}(x)=d_{M}(x,\Sigma_{p})=d_{M}(x,p_{0})-r,\quad x\in\Omega_{R}.$
Since $\nabla d$ is orthogonal to the level sets $\Sigma_{p,s}:=d^{-1}(s)$, for $s=d(x)$, with $\Sigma_{p}=\Sigma_{0}$ and $s\in[0,R-r)$, we can define an orthonormal frame in a neighborhood of any given point $x\in\Omega_{R}$ as follows. First, near $x\in\Sigma_{p,s}\cap\Omega_{R}$, we choose an orthonormal frame $\{E_{i}\}_{i=1}^{n-1}$ tangent to $\Sigma_{p,s}$. We then extend this frame to a neighborhood $\Lambda\subset\Omega_{R}$ of $x$ by parallel transport along the geodesic rays emanating from $p_{0}$, and we continue to denote the extension by $\{E_{i}\}_{i=1}^{n-1}$. Define $E_{n}:=\nabla d$ on $\Lambda$. Then, $\{E_{i}\}_{i=1}^{n}$ forms an orthonormal frame on $\Lambda\subset\Omega_{R}$. [^1] Without loss of generality, we may assume from the outset that $\Lambda=B_{R}(p_{0})\cap\Omega$, i.e., $\Lambda=\Omega_{R}$.
Observe that, since the vector fields $E_{i}$ are parallel along the geodesics issuing from $p_{0}$, we have $\nabla_{E_{n}}E_{i}=0$ on $\Lambda$ for all $i=1,\dots,n$.
Lemma 5
In a neighborhood $\Lambda\subset\Omega_{R}$ as described above, we have:
$\nabla^{2}d(E_{i},E_{n})=0,\quad\text{for }i=1,\dots,n,$
and
$\nabla^{2}d(E_{i},E_{j})=-H_{E_{n}}(E_{i},E_{j}),\quad\text{for }i,j=1,\dots,n-1,$
where
$H_{E_{n}}(E_{i},E_{j}):=\left\langle-(\nabla_{E_{i}}E_{n})^{\top},E_{j}\right\rangle.$
In particular, $\nabla^{2}d(E_{i},E_{i})=-II_{E_{n}}(E_{i})$, where $II$ denotes the second fundamental form of the hypersurfaces $\Sigma_{p,s}$, and
$\Delta d=-\sum_{i=1}^{n-1}II_{E_{n}}(E_{i}).$
Proof. Let $E_{n}=\nabla d$ and $\{E_{1},\dots,E_{n-1}\}$ tangent to the level set $\Sigma_{p,s}$, extended off $\Sigma_{p,s}$ by parallel transport along the radial geodesics from $p_{0}$. Since $|\nabla d|=1$ on the regular set of $d$, we have
$0=\tfrac{1}{2}E_{i}\!\left(|\nabla d|^{2}\right)=\langle\nabla_{E_{i}}\nabla d,\nabla d\rangle=\nabla^{2}d(E_{i},E_{n})$
for every $i=1,\dots,n$, proving the first assertion.
For $i,j\leq n-1$, $\nabla^{2}d(E_{i},E_{j})=\langle\nabla_{E_{i}}E_{n},E_{j}\rangle$. Decompose $\nabla_{E_{i}}E_{n}$ into tangential and normal parts relative to $\Sigma_{p,s}$: $\nabla_{E_{i}}E_{n}=(\nabla_{E_{i}}E_{n})^{\top}+(\nabla_{E_{i}}E_{n})^{\perp}$. But $E_{n}$ is the unit normal to $\Sigma_{p,s}$, thus the Weingarten map $A_{E_{n}}$ is defined by $A_{E_{n}}(E_{i})=-(\nabla_{E_{i}}E_{n})^{\top}$ and the second fundamental form by $II_{E_{n}}(E_{i},E_{j})=\langle A_{E_{n}}(E_{i}),E_{j}\rangle$. Hence
$\nabla^{2}d(E_{i},E_{j})=\langle\nabla_{E_{i}}E_{n},E_{j}\rangle=\left\langle(\nabla_{E_{i}}E_{n})^{\top},E_{j}\right\rangle=-\,II_{E_{n}}(E_{i},E_{j}),$
which yields the stated identity with $H_{E_{n}}(E_{i},E_{j}):=\langle-(\nabla_{E_{i}}E_{n})^{\top},E_{j}\rangle$. Taking the trace over the tangential directions,
$\Delta d=\sum_{k=1}^{n}\nabla^{2}d(E_{k},E_{k})=\sum_{i=1}^{n-1}\nabla^{2}d(E_{i},E_{i})+\nabla^{2}d(E_{n},E_{n})=-\sum_{i=1}^{n-1}II_{E_{n}}(E_{i}),$
since $\nabla^{2}d(E_{n},E_{n})=\langle\nabla_{E_{n}}E_{n},E_{n}\rangle=0$ by $|\nabla d|=1$ and metric-compatibility.
Lemma 6
Let $\varphi\in C^{2}(\overline{\Omega})$ and $\psi\in C^{\infty}([0,\infty))$ be given. Define the function $w\in C^{2}(\overline{\Lambda})$ by
$w=\varphi+\psi\circ d,$ (21)
where $\Lambda\subset\Omega_{R}$ is a neighborhood as defined above. Then,
$\nabla^{2}w(\nabla w,\nabla w)$ $=\left|\nabla\varphi\right|^{2}\nabla^{2}\varphi\left(\frac{\nabla\varphi}{\left|\nabla\varphi\right|},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)$
$\quad+\psi^{\prime}(d)\left[2\left|\nabla\varphi\right|\nabla^{2}\varphi\left(E_{n},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)+\left|\nabla\varphi\right|^{2}\nabla^{2}d\left(\frac{\nabla\varphi}{\left|\nabla\varphi\right|},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)\right]$
$\quad+\left[\psi^{\prime}(d)\right]^{2}\nabla^{2}\varphi(E_{n},E_{n})+\psi^{\prime\prime}(d)\left(E_{n}(\varphi)+\psi^{\prime}(d)\right)^{2}.$ (22)
Proof. We compute $\nabla_{\nabla w}\nabla w$ using linearity and the chain rule. Since $w=\varphi+\psi(d)$ and $\nabla(\psi\circ d)=\psi^{\prime}(d)\nabla d=\psi^{\prime}(d)E_{n}$,
$\nabla_{\nabla w}\nabla w=\nabla_{\nabla\varphi}\nabla\varphi+\psi^{\prime}(d)\left[\nabla_{\nabla\varphi}E_{n}+\nabla_{E_{n}}\nabla\varphi+\psi^{\prime\prime}(d)E_{n}\right]+\psi^{\prime\prime}(d)E_{n}(\varphi)E_{n}.$
Taking the inner product with $\nabla w=\nabla\varphi+\psi^{\prime}(d)E_{n}$ and noting $\langle\nabla_{\nabla\varphi}E_{n},E_{n}\rangle=\nolinebreak 0$ (since $\langle E_{n},E_{n}\rangle=1$), we obtain
$\langle\nabla_{\nabla w}\nabla w,\nabla w\rangle$ $=\nabla^{2}\varphi(\nabla\varphi,\nabla\varphi)+\psi^{\prime}(d)\big[2\,\nabla^{2}\varphi(E_{n},\nabla\varphi)+\nabla^{2}d(\nabla\varphi,\nabla\varphi)\big]$
$\quad+\big[\psi^{\prime}(d)\big]^{2}\nabla^{2}\varphi(E_{n},E_{n})+\psi^{\prime\prime}(d)\big(E_{n}(\varphi)+\psi^{\prime}(d)\big)^{2},$
which is (22) after factoring $|\nabla\varphi|$ where appropriate.
Proposition 7
Assume that $\partial\Omega^{-}$ satisfies the locally Hadamard $r$-es condition. Let $\varphi\in C^{2}\left(\overline{\Omega}\right)$ be given, and let $\sigma:=\varrho\left(1-\theta\right)$, where $\varrho$ and $\theta$ are given by (12) and (16), respectively. Consider the function
$\psi(s)=\frac{1}{\sigma}\left[\varrho s+\left(\frac{\sigma-\varrho}{\sigma}\right)\left(e^{\sigma s}-1\right)\right],\quad s\in[0,\delta],$ (23)
where
$\delta<\min\left\{R-r,\frac{1}{\sigma}\ln\left(\frac{\varrho}{\varrho-\sigma}\right)\right\},$
and set $\Lambda_{p,\delta}=B_{r+\delta}(p_{0})\cap\Omega$. Then the function $w=\varphi+\psi\circ d$ belongs to $C^{2}(\overline{\Lambda}_{p,\delta})$ and satisfies
$w(p)=\varphi(p),\quad w(x)\geq\varphi(x),\quad\text{and}\quad\mathcal{M}(w(x))\leq 0\quad\text{for all }x\in\overline{\Lambda}_{p,\delta}.$
Proof. We explain how the function (23) was constructed. We start by assuming only that $\psi\in C^{\infty}\left([0,\infty\right)$ and satisfies the conditions $\psi\left(0\right)=0$, $\psi^{\prime}\left(s\right)>0$ and $\psi^{\prime\prime}\left(s\right)<0$ for $s>0$, necessary conditions for $\psi$ to serve as an upper barrier. We assume, from now onward, that $\Lambda\subset\Omega_{R}$ is such that $p\in\partial\Lambda$. From (1), we have $\mathcal{M}(w)(x)\leq 0$ in $\Lambda$ if
$(1+|\nabla w|^{2})\Delta w(x)-\nabla^{2}w(\nabla w,\nabla w)(x)\leq 0,$
where $w=\varphi+\psi\circ d$. By Lemma 6, $\nabla^{2}w(\nabla w,\nabla w)$ can be expressed as in (22). Since
$\left(1+\left|\nabla w\right|^{2}\right)\Delta w=\left(1+\left|\nabla w\right|^{2}\right)\Delta\varphi+(1+\left|\nabla w\right|^{2})\psi^{\prime\prime}\left(d\right)+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d,$
it follows that $\mathcal{M}(w)(x)\leq 0$ if the following sum is nonpositive:
$\left(1+\left|\nabla w\right|^{2}\right)\Delta\varphi+(1+\left|\nabla w\right|^{2})\psi^{\prime\prime}\left(d\right)+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d+$
$-\left|\nabla\varphi\right|^{2}\nabla^{2}\varphi\left(\frac{\nabla\varphi}{\left|\nabla\varphi\right|},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)+$
$-2\psi^{\prime}\left(d\right)\left|\nabla\varphi\right|\nabla^{2}\varphi\left(E_{n},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)-\psi^{\prime}\left(d\right)\left|\nabla\varphi\right|^{2}\nabla^{2}d\left(\frac{\nabla\varphi}{\left|\nabla\varphi\right|},\frac{\nabla\varphi}{\left|\nabla\varphi\right|}\right)+$
$-\left[\psi^{\prime}\left(d\right)\right]^{2}\nabla^{2}\varphi(E_{n},E_{n})-\psi^{\prime\prime}\left(d\right)\left(E_{n}\left(\varphi\right)+\psi^{\prime}\left(d\right)\right)^{2}.$
In $\Lambda$, we estimate:
$-|\nabla\varphi|^{2}\nabla^{2}\varphi\left(\frac{\nabla\varphi}{|\nabla\varphi|},\frac{\nabla\varphi}{|\nabla\varphi|}\right)$ $\leq\tau^{3},$
$-|\nabla\varphi|\nabla^{2}\varphi\left(E_{n},\frac{\nabla\varphi}{|\nabla\varphi|}\right)$ $\leq\tau^{2},$
$-|\nabla\varphi|^{2}\nabla^{2}d\left(\frac{\nabla\varphi}{|\nabla\varphi|},\frac{\nabla\varphi}{|\nabla\varphi|}\right)$ $\leq\tau^{2}\mu_{r},$
$-\nabla^{2}\varphi(E_{n},E_{n})$ $\leq\tau.$
where $\tau$ is given by (11). The third inequality uses that $\nabla^{2}d_{p}|_{\Sigma_{p,s}}\leq\mu_{r}$ for all $s\in[0,R]$, due to the locally Hadamard $r$-es condition. Therefore, the sum is nonpositive in $\Lambda$ if
$\left(1+\left|\nabla w\right|^{2}\right)\tau+(1+\left|\nabla w\right|^{2})\psi^{\prime\prime}\left(d\right)+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d+$
$+\tau^{3}+2\psi^{\prime}\left(d\right)\tau^{2}+\psi^{\prime}\left(d\right)\tau^{2}\mu_{r}$
$+\left[\psi^{\prime}\left(d\right)\right]^{2}\tau-\psi^{\prime\prime}\left(d\right)\left(E_{n}\left(\varphi\right)+\psi^{\prime}\left(d\right)\right)^{2}$ (24)
is non positive. Now compute
$\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime\prime}\left(d\right)$ $=$ $\psi^{\prime\prime}\left(d\right)+\left|\nabla\varphi\right|^{2}\psi^{\prime\prime}\left(d\right)+$
$+2\psi^{\prime}\left(d\right)E_{n}\left(\varphi\right)\psi^{\prime\prime}\left(d\right)+\left[\psi^{\prime}\left(d\right)\right]^{2}\psi^{\prime\prime}\left(d\right)$
and, since $\psi^{\prime\prime}<0$, we obtain
$\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime\prime}\left(d\right)-\psi^{\prime\prime}\left(d\right)\left(E_{n}\left(\varphi\right)+\psi^{\prime}\left(d\right)\right)^{2}\leq\psi^{\prime\prime}\left(d\right).$
Then (24) is non positive in $\Lambda$ if
$\left(1+\left|\nabla w\right|^{2}\right)\tau+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d+$
$\tau^{3}+2\psi^{\prime}\left(d\right)\tau^{2}+\psi^{\prime}\left(d\right)\tau^{2}\mu_{r}$
$+\left[\psi^{\prime}\left(d\right)\right]^{2}\tau+\psi^{\prime\prime}\left(d\right)\leq 0.$ (25)
We have
$\left(1+\left|\nabla w\right|^{2}\right)\tau$ $\leq$ $\left(1+\left|\nabla\varphi\right|^{2}+2\psi^{\prime}\left(d\right)\left|\nabla\varphi\right|+\left[\psi^{\prime}\left(d\right)\right]^{2}\right)\tau$
$\leq$ $\tau+\tau^{3}+2\tau^{2}\psi^{\prime}\left(d\right)+\tau\left[\psi^{\prime}\left(d\right)\right]^{2}.$
Thus, inequality (25) holds if
$\tau+2\tau^{3}+4\tau^{2}\psi^{\prime}\left(d\right)+2\tau\left[\psi^{\prime}\left(d\right)\right]^{2}+\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\Delta d+$
$+\psi^{\prime}\left(d\right)\tau^{2}\mu_{r}+\psi^{\prime\prime}\left(d\right)$ (26)
is nonpositive. Using Lemma 5, for $x\in\Lambda$ and $s=d(x)\in[0,R)$, we have
$\Delta d(x)=-\sum_{i=1}^{n-1}II_{E_{n}}(E_{i})(x)=-(n-1)H_{\Sigma_{p,s}}(x),$
where $H_{\Sigma_{p,s}}$ is the mean curvature of the immersion $\Sigma_{p,s}$ with respect to $E_{n}=\nabla d$. Since $p\in\partial\Omega^{-}$, we have
$H_{\Sigma_{p}}(p)\leq H_{\partial\Omega}(p)<0,$
and from the locally Hadamard $r$-es condition,
$\lambda_{r}\leq H_{\Sigma_{p}}(p)\leq H_{\Sigma_{p,s}}(x),$
for all $x\in\Sigma_{p,s}\cap\Omega_{R}$ and $s\in[0,R-r]$. Hence,
$\Delta d(x)\leq-(n-1)\lambda_{r}$
and we have (26) if
$\tau+2\tau^{3}+4\tau^{2}\psi^{\prime}\left(d\right)+2\tau\left[\psi^{\prime}\left(d\right)\right]^{2}-\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\left(n-1\right)\lambda_{r}+$
$+\psi^{\prime}\left(d\right)\tau^{2}\mu_{r}+\psi^{\prime\prime}\left(d\right)\leq 0.$ (27)
Observing that
$\left(1+\left|\nabla w\right|^{2}\right)\psi^{\prime}\left(d\right)\leq\left(1+\tau^{2}\right)\psi^{\prime}\left(d\right)+2\left[\psi^{\prime}\left(d\right)\right]^{2}\tau\mathfrak{+}\left[\psi^{\prime}\left(d\right)\right]^{3}$
it follows that we have (27) in $\Lambda$ if
$\psi^{\prime\prime}\left(d\right)$ $\leq$ $\left(n-1\right)\lambda_{r}\left[\psi^{\prime}\left(d\right)\right]^{3}-2\tau\left[1-\left(n-1\right)\lambda_{r}\right]\left[\psi^{\prime}\left(d\right)\right]^{2}$
$-\left[\left(4+\mu_{r}\right)\tau^{2}-\left(1+\tau^{2}\right)\left(n-1\right)\lambda_{r}\right]\psi^{\prime}\left(d\right)$
$-\tau\left[1+2\tau^{2}\right].$
Letting
$\alpha:=2\tau[1-(n-1)\lambda_{r}],\quad\beta:=(4+\mu_{r})\tau^{2}-(1+\tau^{2})(n-1)\lambda_{r},$
we write the inequality above as
$\psi^{\prime\prime}\leq\left(n-1\right)\lambda_{r}\left[\psi^{\prime}\right]^{3}-\alpha\left[\psi^{\prime}\right]^{2}-\beta\psi^{\prime}-\varrho.$ (28)
Assume $0<\psi^{\prime}\leq 1$. Then
$\left(n-1\right)\lambda_{r}\psi^{\prime}$ $\leq$ $\left(n-1\right)\lambda_{r}\left[\psi^{\prime}\right]^{3}<0$
$-\alpha\psi^{\prime}$ $\leq$ $-\alpha\left[\psi^{\prime}\right]^{2}<0$
so inequality (28) is satisfied if $\psi^{\prime\prime}-\sigma\psi^{\prime}+\varrho\leq 0$, with $\sigma=\varrho(1-\theta)<0$ since $\theta>1$. Thus, we solve the ODE
$\psi^{\prime\prime}-\sigma\psi^{\prime}+\varrho=0,$ (29)
with initial condition $\psi(0)=0$, whose solution is
$\psi(s)=\frac{1}{\sigma}\left[\varrho s-c_{1}(e^{\sigma s}-1)\right].$
We require $0<\psi^{\prime}(s)\leq 1$, i.e.,
$0<\frac{1}{\sigma}(\varrho-c_{1}\sigma e^{\sigma s})\leq 1,$
which implies
$\frac{\varrho}{\sigma}e^{-\sigma s}<-c_{1}\leq\left(\frac{\sigma-\varrho}{\sigma}\right)e^{-\sigma s}.$
Let $\rho=R-r$. Then everything is in order if
$\frac{\varrho-\sigma}{\sigma}\leq c_{1}<\frac{\varrho}{\sigma}e^{-\sigma\rho}.$ (30)
This holds if
$\ln\left(\frac{\varrho}{\varrho-\sigma}\right)<\sigma\rho.$ (31)
As $\frac{\varrho}{\varrho-\sigma}<1$, condition (31) is satisfied if
$\rho<\frac{1}{\sigma}\ln\left(\frac{\varrho}{\varrho-\sigma}\right).$
Therefore, if
$\delta<\min\left\{R-r,\frac{1}{\sigma}\ln\left(\frac{\varrho}{\varrho-\sigma}\right)\right\},$ (32)
then, for
$\psi(s)=\frac{1}{\sigma}\left[\varrho s-c_{1}(e^{\sigma s}-1)\right],\quad s\in[0,\delta],$
with $c_{1}=\frac{\varrho-\sigma}{\sigma}$, we have $\psi(0)=0$, $0<\psi^{\prime}(s)\leq 1$, and $\psi^{\prime\prime}(s)=-c_{1}\sigma e^{\sigma s}<0$. Furthermore, $\psi(s)\geq 0$ on $0,\delta]$. Finally, defining
$\Lambda_{p,\delta}=B_{r+\delta}(p_{0})\cap\Omega,$
we obtain the function
$w(x)=\varphi(x)+\psi(d(x)),$
with
$\psi(s)=\frac{1}{\sigma}\left[\varrho s+\left(\frac{\sigma-\varrho}{\sigma}\right)(e^{\sigma s}-1)\right],\quad 0\leq s\leq\delta,$ (33)
which satisfies $w(p)=\varphi(p)$, $w(x)\geq\varphi(x)$ for all $x\in\overline{\Lambda}_{p,\delta}$, and $\mathcal{M}(w(x))\leq 0$ in $\overline{\Lambda}_{p,\delta}$.
1.2 Proofs of Theorem 3 and Corollary 4
Proof of Theorem 3.
In the case where $\partial\Omega^{-}=\emptyset$ and $\Omega$ is bounded, the result follows, for instance, from Theorem 1 of [3].
Suppose $\Omega$ bounded and that $\partial\Omega^{-}$ satisfies the locally Hadamard $r$-es condition.
Given any $p\in\partial\Omega^{-}$, consider the neighborhood $\Lambda_{p,\delta}$ as defined in Proposition 7, where $\delta$ satisfies condition (32). Define the function
$w_{p}=\varphi+\psi(d_{p}),$
where $\psi$ is as in (33). To construct a barrier from above at $p$, we require that
$w_{p}(x)=\varphi(x)+\psi(\delta)\geq\sup_{\overline{\Omega}}\varphi\quad\text{on }\partial\Lambda_{p,\delta}\setminus\partial\Omega,$
which holds if
$\psi(\delta)\geq\sup_{\overline{\Omega}}\varphi-\inf_{\overline{\Omega}}\varphi=:\omega.$
Since $\delta$ is independent of the point $p\in\partial\Omega^{-}$, Proposition 7 implies that, if
$\omega\leq\frac{1}{\sigma}\left[\varrho\delta+\left(\frac{\sigma-\varrho}{\sigma}\right)\left(e^{\sigma\delta}-1\right)\right],$ (34)
(which is the expression (17) since that $\sigma=\varrho(1-\theta)$), then we can construct an upper barrier at each $p\in\partial\Omega^{-}$ by setting
$v_{p}(x)=\begin{cases}\min\left\{w_{p}(x),\inf_{\partial\Lambda_{p,\delta}\setminus\partial\Omega}w_{p}\right\}&\text{if }x\in\Lambda_{p,\delta},\\ \inf_{\partial\Lambda_{p,\delta}\setminus\partial\Omega}w_{p},&\text{if }x\in\Omega\setminus\Lambda_{p,\delta}.\end{cases}$
For a lower barrier at $p\in\partial\Omega^{-}$, just work with $\xi_{p}:=\varphi-\psi\circ d$.
At points of $\partial\Omega$ where the domain is mean convex, we proceed as in the proof of Theorem 1 of [3], and no additional geometric assumptions are required to construct barriers.
The conclusion now follows from classical elliptic PDE theory [10].
If $\Omega$ is unbounded and $\tau<\infty$, the result follows directly from the method of Perron (see Theorem 5 of [15]). Indeed, define, at given $x\in\overline{\Omega},$
$u\left(x\right)=\sup\left\{s\left(x\right);s\in C^{0}\left(\Omega\right)\text{ is generalized subsolution of }\mathcal{M}\text{ and }s|_{\partial\Omega}\leq\varphi|_{\partial\Omega}\right\}.$
Since every subsolution used to define $u$ is bounded above by $\sup_{\Omega}\varphi$, the function $u$ is well-defined. Moreover, by Perron’s method, we have $u\in\nolinebreak C^{\infty}\left(\Omega\right)$ and $\mathcal{M}\left(u\right)=0.$ The barriers at the boundary ensures that $u\in C^{2}\left(\overline{\Omega}\right)$ and $u|_{\partial\Omega}=\varphi.$
Proof of Corollary 4.
When $M$ is a Hadamard manifold, we have $R=\infty$, so we can take $\delta$ in (32) as
$\delta=\frac{1}{\sigma}\ln\left(\frac{\varrho}{\varrho-\sigma}\right).$
Substituting this into (34) and recalling that $\sigma=\varrho(1-\theta)$, condition (34) becomes
$\omega\leq\frac{1}{\varrho(\theta-1)^{2}}\left(\theta-\ln\theta-1\right).$ (35)
Since $1<a<\theta$ and observing that the function $f(x)=x-\ln x-1$ is positive and increasing for $x>1$, we have that
$\omega\leq\frac{a-\ln a-1}{\varrho(\theta-1)^{2}}$ (36)
implies (35). Now observe that
$\theta=a+b(n-1)\lambda_{r}+c\mu_{r}.$
Plugging this into (36) leads to the equivalent inequality
$\frac{\sqrt{a-\ln a-1}-\sqrt{\varrho\omega}(a+c\mu_{r}-1)}{\sqrt{\varrho\omega}\,b(n-1)}\leq\lambda_{r}.$ (37)
Remark 8
Since $\lambda_{r}<0$, the left-hand side of (37) must be negative. This occurs if
$\omega<\frac{a-\ln a-1}{\varrho(a+c\mu_{r}-1)^{2}},$
which is guaranteed by condition (36).
Example 9
We now present an example illustrating that our condition ensures the existence of a solution to the Dirichlet problem (2) even in cases where the Jenkins-Serrin criterion fails to do so.
Let $\Omega\subset\mathbb{R}^{2}$ be the bounded domain whose boundary is parametrized by
$\gamma(t)=\xi(t)(\cos t,\sin t),\quad t\in[0,2\pi],$
where
$\xi(t)=\sqrt{4\cos(2t)+\sqrt{16\cos^{2}(2t)+12}}.$
Let $\varphi\in C^{2}(\overline{\Omega})$ be given by
$\varphi(x,y)=\frac{e^{-y}}{20}+1.$
Figure 1: The domain used in the example
Regarding the geometry of $\partial\Omega$, we compute
$0.45\approx\lambda_{r}\leq k_{\partial\Omega}\leq k:=\sup_{\partial\Omega}|k_{\partial\Omega}|\approx 0.82,$
so that $\mu_{r}=-\lambda_{r}$. For the boundary data $\varphi|_{\partial\Omega}$, we find
$\omega_{\varphi}(\overline{\Omega})=\omega_{\varphi}(\partial\Omega)\approx 0.17,$
with $\tau\approx 0.1871$, $\theta\approx 9.05$, and the maximal admissible value of $\delta$ from (32) is approximately $1.37$. Choosing $\delta=1$ for simplicity, the right-hand side of inequality (17) becomes approximately $0.43$. Since $\omega\approx 0.17<0.43$, the condition is satisfied.
On the other hand, following the Jenkins-Serrin method, we proceed as in case I) above for computing $B$. Using the corresponding notation and in order to obtain the maximum value for $B$, we estimate:
$l=0.25,\quad l^{\prime}\approx 0.1767,\quad\frac{\pi}{l^{\prime}}\approx 17.7792,\quad k\approx 0.82,$
$|D^{2}\varphi|\approx 0.1871,\quad A\approx 17.7792,\quad C\approx 10.2526,\quad\mathcal{H}\approx 0.45.$
These values yield
$B\approx 0.0072<\omega_{\varphi}(\partial\Omega),$
so the Jenkins-Serrin condition fails, while our condition still guarantees existence.
2 Mean Curvature Flow of Graphs without Mean Convexity
The differential inequality leading to the ordinary differential equation
$\psi^{\prime\prime}-\sigma\,\psi^{\prime}+\varrho=0,$ (38)
introduced in the previous section, naturally reappears in the study of the mean curvature flow of graphs. The same function $\psi$ that provides elliptic barriers for the Dirichlet problem also yields parabolic barriers that allow one to start and control the evolution even when $\partial\Omega$ is not mean convex. Thus equation (38) plays a central role in connecting the elliptic and parabolic regimes.
In the elliptic construction, the auxiliary function is written in terms of the distance $d_{p}(x)=\mathrm{dist}(x,\Sigma_{p})$ to the exterior geodesic sphere $\Sigma_{p}$ associated with each point $p\in\partial\Omega^{-}$, since these spheres encode the geometric rr–exterior sphere condition.
In the parabolic setting, in order to obtain barriers that are compatible with the Dirichlet condition along the whole boundary, it is more natural to work with the distance to the boundary,
$d(x)=\mathrm{dist}(x,\partial\Omega).$
Although $d(x)$ and $d_{p}(x)$ need not to coincide as functions in a neighborhood of $p$, along $\partial\Omega$ the two functions have the same first and second derivatives in the normal direction at $p$, that is, their first and second jets at $p$ coincide. Since all the differential inequalities leading to the ODE (38) depend only on the values of $\nabla d$ and $\nabla^{2}d$ along the boundary, this second–order agreement is sufficient to transfer the elliptic barrier construction to the parabolic framework without any modification of the resulting ODE.
Let $\Omega\subset M^{n}$ be a $C^{2}$ domain in a complete Riemannian manifold and let $\varphi\in C^{2}(\overline{\Omega})$ satisfy the structural condition of Theorem 3 (or, in the Hadamard case, condition (19)). We conclude from the above arguments that there is a tubular neighborhood of $\partial\Omega$,
$U_{\delta}:=\{x\in\Omega:d(x)<\delta\},$
where the signed distance function $d$ to $\partial\Omega$ is of class $C^{2}$ and, chosen $\psi$ as in Section 1.1 (Proposition 7),
$w^{\pm}(x):=\varphi(x)\pm\psi(d(x)),\qquad x\in U_{\delta},$
are such that $w^{-}$ is a subsolution and $w^{+}$ is a supersolution for the operator $\mathcal{M}$ in $U_{\delta}$.
Theorem 10
Let $\Omega\subset M^{n}$ be a $C^{2}$ domain in a complete Riemannian manifold, and let $\varphi\in C^{2}(\overline{\Omega})$ satisfy the structural condition of Theorem 3 (or, in the Hadamard case, condition (19)). Let $u_{0}\in C^{2}(\overline{\Omega})$ satisfy
$u_{0}|_{\partial\Omega}=\varphi,~w^{-}(x)\leq u_{0}(x)\leq w^{+}(x)\text{ for all }x\in U_{\delta},\text{ }\left|u_{0}\right|\leq\inf_{\partial U_{\delta}\backslash\partial\Omega}w^{+}\text{,}$
where $w^{\pm}$ and $U_{\delta}$ are as defined above. Then there exists $T>0$, depending only on $\varphi$ (through $\tau(\varphi)$) and on the geometric parameters $(\lambda_{r},\mu_{r})$ arising in the construction of the boundary barriers and, in particular, independent of the choice of $u_{0}$, such that the initial–boundary value problem
$\begin{cases}\partial_{t}u=\mathcal{M}(u),&(x,t)\in\Omega\times(0,T),\\ u(x,0)=u_{0}(x),&x\in\Omega,\\ u(x,t)=\varphi(x),&(x,t)\in\partial\Omega\times[0,T),\end{cases}$ (39)
admits a classical solution
$u\in C^{2,1}(\overline{\Omega}\times[0,T]).$
Moreover, for every $t\in[0,T]$, the hypersurface
$\Gamma_{t}=\{(x,u(x,t)):x\in\Omega\}$
remains a graph over $\Omega$, even if $\partial\Omega$ is not mean convex.
Proof. The proof relies directly on the equation (38). Let $d(x)$ denote the signed distance to $\partial\Omega$ and consider $w^{\pm}$ and $U_{\delta}$ are as defined above.
Since $w^{\pm}$ are independent of $t$, we have $\partial_{t}w^{\pm}=0$, and because $\psi$ satisfies (38), the functions $w^{+}$ and $w^{-}$ verify
$\partial_{t}w^{+}-\mathcal{M}(w^{+})\geq 0,\qquad\partial_{t}w^{-}-\mathcal{M}(w^{-})\leq 0$
in the tubular neighborhood $U_{\delta}$ of $\partial\Omega$, and both coincide with $\varphi$ on $\partial\Omega$.
Under the quantitative restriction (17) (or (19)), we have $w^{-}(x)\leq\varphi(x)\leq w^{+}(x)$ on $\partial\Omega$ and $w^{-}\leq w^{+}$ throughout $U_{\delta}$. To apply the parabolic maximum principle globally, we extend $w^{\pm}$ from $U_{\delta}$ to whole $\Omega$ by a standard $C^{2}$ gluing argument, by smoothing the capped function
$W^{+}\left(x\right)=\left\{\begin{array}[]{c}\min\{w^{+}\left(x\right),\inf_{\partial\Omega_{\delta}\backslash\partial\Omega}w^{+}\}\text{ se }x\in\overline{\Omega}_{\delta}\\ \inf_{\partial\Omega_{\delta}\backslash\partial\Omega}w^{+}\text{ se }x\in\Omega\backslash\Omega_{\delta}\end{array}\right$.
(which is a $C^{0}$ supersolution in $\overline{\Omega}$ for the operator $\mathcal{M}$) in such a way that the supersolution property is preserved, and we call it again $W^{+}$ (analogous construction for $W^{-}$). From our hypothesis on $u_{0}$, it follows that $W^{-}\leq u_{0}\leq W^{+}$ on $\overline{\Omega}$ with $u_{0}=\varphi$ on $\partial\Omega$. The classical theory for quasilinear parabolic equations (see [13]) guarantees local existence and uniqueness provided the structure is uniformly parabolic. We thus verify this uniform parabolicity within the region bounded by $W^{-}$ and $W^{+}$.
Since $\psi$ is smooth and monotone, the derivatives $|\nabla w^{\pm}|$ are bounded by constants depending only on $\tau(\varphi)$ and $(\lambda_{r},\mu_{r})$. By the parabolic maximum principle, the solution $u$ of (39) remains confined between $W^{-}$ and $W^{+}$, so that
$W^{-}(x)\leq u(x,t)\leq W^{+}(x)\quad\text{for all }(x,t)\in\overline{\Omega}\times[0,T].$
Hence $|\nabla u|$ is uniformly bounded because $\psi^{\prime}(d)$ is bounded, and therefore the coefficients
$a^{ij}(\nabla u)=\delta^{ij}(1+|\nabla u|^{2})-u_{i}u_{j}$
remain positive definite. The equation (39) is thus uniformly parabolic during the time interval considered.
With this structure established, the fixed–point argument applies: one linearizes (39) around $u_{0}$, solves the linearized problem via the Schauder theory, and iterates. The limit yields a solution $u\in C^{2,1}$ on $[0,T]$, where $T$ depends only on the upper bound of $|\nabla u|$ obtained from the barriers, hence only on $\tau(\varphi)$, $\lambda_{r}$ and $\mu_{r}$.
The boundary control provided by $\psi$ replaces the classical mean convexity assumption. Since the gradient is uniformly bounded, the graph cannot develop vertical tangents and remains graphical for all $t\in[0,T]$. Interior gradient and curvature estimates, identical to those derived by Ecker–Huisken [6], ensure that no curvature blow–up occurs before $T$. Finally, $u(x,t)\rightarrow u_{0}(x)$ as $t\rightarrow 0$ by parabolic regularity and compatibility of the initial and boundary data.
Consequently, the function $\psi$ solving the ODE (38) provides a universal stationary parabolic barrier that replaces the mean convexity of the boundary, establishing local–in–time existence of the mean curvature flow of graphs over arbitrary domains.
Remark 11
Let $[0,U)$ denote the maximal time interval of existence of the classical solution $u$ to (39). By the interior estimates of Ecker–Huisken, for any sequence $t_{k}\nearrow U$ there exists a subsequence (still denoted $t_{k}$) such that $u(\cdot,t_{k})$ converges in $C_{\mathrm{loc}}^{\infty}(\Omega)$ to a function $v\in C^{\infty}(\Omega)$. Moreover, $v$ satisfies the stationary minimal surface equation $\mathcal{M}(v)=0$ in $\Omega$, hence defines a minimal graph in the interior of the domain. However, since all these estimates are purely interior, no information can be extracted in general about the behavior of the solution at the boundary as $t\nearrow U$. In particular, the possible obstruction to extending the flow beyond $U$ is entirely of boundary nature.